# Localized RETE for Incremental Graph Queries with Nested Graph Conditions

### v1

@@ 4.2. Localized Detection of NGC Satisfaction Changes
While RETE nets constructed via $localize^{\Psi}$ can be used to locally and incrementally compute matches for an extended query that touch a relevant subgraph, the notion of completeness in Definition 4.1 no longer covers the use case of correctly tracking the full effect of local changes on a global query result. This is due to the fact that for extended queries, local changes may result in changes to the satisfaction a nested graph condition for matches that do not touch the relevant subgraph, effectively manipulating query results that do not fall under Definition 4.1.
Notably, nested graph conditions have been shown to be equivalent to first-order logic over graphs [HP09, Ren04]. The satisfiability problem for first-order logic being undecidable thus makes a direct computation of matches for which NGC satisfaction somehow depends on the relevant subgraph impossible in the general case. In order to enable a localized tracking of changes to the results of an extended graph query, we therefore employ the following overapproximation:
@@
  • $\psi=\exists(a:Q\rightarrow Q^{\prime},\psi^{\prime})$ and there exists a graph morphism $m^{\prime}:Q^{\prime}\rightarrow H$ such that $m=m^{\prime}\circ a$ and $m^{\prime}(Q^{\prime})\cap H_{p}\neq\emptyset\vee m^{\prime}\text{ is subgraphsatisfaction dependent}$
@@
We furthermore define a subgraph-restricted, that is, local graph modification as a graph modification where changes are limited to relevant subgraphs of source and target graph and relevant subgraphs are mapped onto each other:
@@
  • For a nested graph condition of the form $\psi=\exists(a:Q\rightarrow Q^{\prime},\psi^{\prime})$, $(N^{sat},p^{sat})=localize^{sat}(Q,\psi)$ consists of the localized RETE net for the plain pattern $Q$, $(N^{\Phi}_{Q},p^{\Phi}_{Q})=localize(Q)$, the localized RETE net for the plain pattern $Q^{\prime}$, $(N^{\Phi}_{Q^{\prime}},p^{\Phi}_{Q^{\prime}})=localize(Q^{\prime})$, the RETE net $(N^{sat}_{(Q^{\prime},\psi^{\prime})},p^{sat}_{(Q^{\prime},\psi^{\prime})})=localize^{sat}(Q^{\prime},\psi^{\prime})$, a marking-sensitive union node $[\cup]^{\Phi}$ with dependencies $p^{\Phi}_{Q^{\prime}}$ and $p^{sat}_{(Q^{\prime},\psi^{\prime})}$, a request projection structure $RPS_{r}=RPS([\cup]^{\Phi},N^{\Phi}_{Q})$, and a marking-sensitive semi-join node $p^{sat}=[\ltimes]^{\Phi}$ along $a$ with left dependency $p^{\Phi}_{Q}$ and right dependency $[\cup]^{\Phi}$.
@@
A consistent configuration for the resulting RETE net then guarantees a complete query result with respect to Definition 4.2:
@@
Let $H$ be a graph, $H_{p}\subseteq H$, and $(Q,\psi)$ an extended graph query. Furthermore, let $\mathcal{C}^{\Phi}$ be a consistent configuration for the RETE net $(N^{sat},p^{sat})=localize^{sat}(Q,\psi)$. It then holds that $\forall m\in\mathcal{M}^{Q}_{H}:m\text{ is subgraph satisfaction dependent }\Rightarrow m\in\mathcal{R}^{\Phi}_{X}(p^{sat},\mathcal{C}^{\Phi})$.
@@
  • For a nested graph condition of the form $\psi=\exists(a:Q\rightarrow Q^{\prime},\psi^{\prime})$, the result of $order^{sat}$ is given by $order^{sat}(N^{sat})=order(N^{\Phi}_{Q^{\prime}})\circ order^{sat}(N^{sat}_{(Q^{\prime},\psi^{\prime})})\circ[\cup]^{\Phi}\circ toposort(RPS_{r})^{-1}\circ order(N^{\Phi}_{Q})\circ[\ltimes]^{\Phi}$.
@@
Let $H$ be a graph, $H_{p}\subseteq H$, $(Q,\psi)$ a graph query, and $\mathcal{C}^{\Phi}_{0}$ an arbitrary starting configuration for the marking-sensitive RETE net $(N^{sat},p^{sat})=localize^{sat}(Q,\psi)$. Executing $(N^{sat},p^{sat})$ via $O=order^{sat}(N^{sat})$ then yields a consistent configuration $\mathcal{C}^{\Phi}=execute(O,N^{sat},H,H_{p},\mathcal{C}^{\Phi}_{0})$.
@@
Let $H$ be a graph, $H_{p}\subseteq H$, and $(Q,\psi)$ a graph query. Furthermore, let $\mathcal{C}^{\Phi}_{0}$ be an arbitrary starting configuration for the marking-sensitive RETE net $(N^{sat},p^{sat})=localize^{sat}(Q,\psi)$ and $\mathcal{C}^{\Phi}=execute(order^{sat}(N^{sat}),N^{sat},H,H_{p},\mathcal{C}^{\Phi}_{0})$. It then holds that $\forall m\in\mathcal{M}^{Q}_{H}:m\text{ is subgraph satisfaction dependent }\Rightarrow m\in\mathcal{R}^{\Phi}_{X}(p^{sat},\mathcal{C}^{\Phi})$.
@@
The result $(N^{\Delta},p^{\Delta})=localize^{\Delta}(Q,\psi)$ of this construction via the procedure $localize^{\Delta}$ is visualized in Figure 6. The resulting RETE net consists of the localized version $(N^{\Phi}_{Q},p^{\Phi}_{Q})=localize(N^{Q},p^{Q})$ of a regular RETE net $(N^{Q},p^{Q})$ with height $h$ for the plain graph query $Q$, a marking filter node $[\phi>h]^{\Phi}$ with dependency $p^{\Phi}_{Q}$, the RETE net $(N^{sat}_{(Q,\psi)},p^{sat}_{(Q,\psi)})=localize^{sat}(Q,\psi)$, and a marking-sensitive union node $[\cup]^{\Phi}$ with dependencies $[\phi>h]^{\Phi}$ and $p^{sat}_{(Q,\psi)}$. In addition, $N^{\Delta}$ comprises the RETE net $(N^{\Phi}_{(Q,\psi)},p^{\Phi}_{(Q,\psi)})=localize^{\Psi}(Q,\psi)$, a request projection structure $RPS_{l}=RPS([\cup]^{\Phi},N^{\Phi}_{(Q,\psi)})$, and a marking-sensitive semi-join $[\ltimes]^{\Phi}$ with left dependency $[\cup]^{\Phi}$ and right dependency $p^{\Phi}_{(Q,\psi)}$.
Additionally, another copy of the RETE net for computing subgraph satisfaction dependent matches, $(N^{sat^{\prime}}_{(Q,\psi)},p^{sat^{\prime}}_{(Q,\psi)})=localize^{sat}(Q,\psi)$, is created for computing subgraph satisfaction dependent matches in $H^{\prime}$. These matches then have to be translated into matches into $H$ along the graph morphisms $f$ and $g$ via a dedicated marking-sensitive translation node $[\xrightarrow{f\circ g^{-1}}]^{\Phi}$ with dependency $p^{sat^{\prime}}_{(Q,\psi)}$, which is also added to $N^{\Delta}$. The target result set of this marking-sensitve translation node is given by $\mathcal{R}^{\Phi}([\xrightarrow{f\circ g^{-1}}]^{\Phi},N^{\Delta},H,H_{p},\mathcal{C}^{\Phi})=\{(m,\phi)|m\in\mathcal{M}^{Q}_{H}\wedge\exists(m^{\prime},\phi)\in\mathcal{C}^{\Phi}(p^{sat^{\prime}}):f\circ g^{-1}\circ m^{\prime}=m\}$. The node $[\xrightarrow{f\circ g^{-1}}]^{\Phi}$ can then be added as an additional dependency to $[\cup]^{\Phi}$.
@@
An execution order for a RETE net $(N^{\Delta},p^{\Delta})=localize^{\Delta}(Q,\psi)$ can simply be constructed as $order^{\Delta}(N^{\Delta})=order(N^{\Phi}_{Q})\circ[\phi>h]^{\Phi}\circ order^{sat}(N^{sat}_{(Q,\psi)})\circ[\xrightarrow{f\circ g^{-1}}]^{\Phi}\circ[\cup]^{\Phi}\circ toposort(RPS_{l})^{-1}\circ order^{\Psi}(N^{\Phi}_{(Q,\psi)})\circ[\ltimes]^{\Phi}$:
@@
Let $(Q,\psi)$ be an extended graph query and $H$ a host graph with relevant subgraph $H_{p}$ and and $\mathcal{C}^{\Phi}_{0}$ an arbitrary starting configuration for the marking-sensitive RETE net $(N^{\Delta},p^{\Delta})=localize^{\Delta}(Q,\psi)$. Executing $(N^{\Delta},p^{\Delta})$ via $O=order^{\Delta}(N^{\Delta})$ then yields a consistent configuration $\mathcal{C}^{\Phi}=execute(O,N^{\Delta},H,H_{p},\mathcal{C}^{\Phi}_{0})$.
@@ 5. Evaluation
We aim to investigate whether RETE net localization can improve performance of query execution in scenarios where the relevant subgraph constitutes only a fraction of the full model, considering initial query execution time, execution time for incrementally processing model updates, and memory consumption as performance indicators. We experiment [^7] Experiments were executed on a Linux SMP Debian 4.19.67-2 computer with Intel Xeon E5-2630 CPU (2.3 GHz clock rate) and 386 GB system memory running OpenJDK 11.0.6. Reported measurements correspond to the arithmetic mean of measurements for 10 runs. Memory measurements were obtained using the Java Runtime class. To represent graph data, all experiments use EMF [emf] object graphs that enable reverse navigation of edges. Our implementation is available under [BG24c]. More details and query visualizations can be found in [BG24b] and [BG24c]. with the following querying techniques:
@@ Appendix A Technical Details
Theorem 22.
Let $H$ be a graph, $H_{p}\subseteq H$, $(N,p)$ a well-formed RETE net, and $Q=query(p)$. Furthermore, let $\mathcal{C}^{\Phi}$ be a consistent configuration for the localized RETE net $(N^{\Phi},p^{\Phi})=localize(N,p)$. It then holds that $\forall(m,\phi)\in\mathcal{C}^{\Phi}(p^{\Phi}):m(Q)\cap H_{p}\neq\emptyset\Leftrightarrow\phi=\infty$.
@@
Theorem 23.
Let $H$ be an edge-dominated graph, $H_{p}\subseteq H$, $(N,p)$ a well-formed RETE net with $Q=query(p)$, $\mathcal{C}$ a consistent configuration for $(N,p)$ for host graph $H$, and $\mathcal{C}^{\Phi}$ a consistent configuration for the marking-sensitive RETE net $(N^{\Phi},p^{\Phi})=localize(N,p)$ for host graph $H$ and relevant subgraph $H_{p}$. It then holds that $\sum_{n^{\Phi}\in V^{N^{\Phi}}}\sum_{(m,\phi)\in\mathcal{C}(n^{\Phi})}|m|\leq 7\cdot|\mathcal{C}|_{e}$.
@@
Theorem 24.
Let $H$ be a graph, $H_{p}\subseteq H$, and $(Q,\psi)$ an extended graph query. Furthermore, let $\mathcal{C}^{\Phi}$ be a consistent configuration for the localized RETE net $(N^{\Phi},p^{\Phi})=localize^{\Psi}(Q,\psi)$. The set of matches given by the stripped result set $\mathcal{R}^{\Phi}_{X}(p^{\Phi},\mathcal{C}^{\Phi})$ is then correct under $H_{p}$.
@@
Theorem 25.
Let $H$ be a graph, $H_{p}\subseteq H$, $(Q,\psi)$ a graph query, and $\mathcal{C}^{\Phi}_{0}$ an arbitrary starting configuration for the marking-sensitive RETE net $(N^{\Phi},p^{\Phi})=localize^{\Psi}(Q,\psi)$. Executing $(N^{\Phi},p^{\Phi})$ via $O=order^{\Psi}(N^{\Phi})$ then yields a consistent configuration $\mathcal{C}^{\Phi}=execute(O,N^{\Phi},H,H_{p},\mathcal{C}^{\Phi}_{0})$.
@@
Theorem 26.
Let $H$ be an edge-dominated graph, $H_{p}\subseteq H$, $(N,p)$ a RETE net created via the procedure described in [BG23] for the extended graph query $(Q,\psi)$, $\mathcal{C}$ a consistent configuration for $(N,p)$ for host graph $H$, and $\mathcal{C}^{\Phi}$ a consistent configuration for the marking-sensitive RETE net $(N^{\Phi},p^{\Phi})=localize^{\Psi}(Q,\psi)$ corresponding to $(N,p)$ for host graph $H$ and relevant subgraph $H_{p}$. It then holds that $\sum_{n^{\Phi}\in V^{N^{\Phi}}}\sum_{(m,\phi)\in\mathcal{C}^{\Phi}(n^{\Phi})}|m|\leq 7\cdot|\mathcal{C}|_{e}$.
@@
For a nested condition of the form $\psi=\exists(a:Q\rightarrow Q^{\prime},\psi^{\prime})$, $N^{\Phi}$ consists of the localized RETE net for $Q$, $(N^{\Phi}_{Q},p^{\Phi}_{Q})=localize(Q)$, the RETE net $(N^{\Phi}_{(Q^{\prime},\psi^{\prime})},p^{\Phi}_{(Q^{\prime},\psi^{\prime})})=localize^{\Psi}(Q^{\prime},\psi^{\prime})$, a request projection structure $RPS_{l}=RPS(p^{\Phi}_{Q},N^{\Phi}_{(Q^{\prime},\psi^{\prime})})$, and a marking-sensitive semi-join node $p^{\Phi}=[\ltimes]^{\Phi}$.

### v2

@@ 4.2. Localized Detection of NGC Satisfaction Changes
While RETE nets constructed via $localize^{\Psi}$ can be used to locally and incrementally compute matches for an extended query that touch a relevant subgraph, the notion of completeness in Definition 4.1 no longer covers the use case of correctly tracking the full effect of local changes on a global query result. This is due to the fact that for extended queries, local changes may result in changes to the satisfaction of a nested graph condition for matches that do not touch the relevant subgraph, effectively manipulating query results that do not fall under Definition 4.1.
To track the effect of local host graph changes on the global set of results for an extended graph query, we thus also have to maintain the set of matches for which a modification of the relevant subgraph may change the nested graph condition’s satisfaction. We ideally also want to precompute this set of matches before the host graph is modified to enable incremental maintenance. We thus have no prior knowledge regarding the modifications we may have to consider except for the fact that they concern the relevant subgraph.
This proves problematic when considering a case where the relevant subgraph constitutes the entire host graph as an example. In this scenario, checking whether an arbitrary modification to the relevant subgraph may change the satisfaction of a nested graph condition for some context match corresponds to checking satisfiability of the nested graph condition.
Notably however, nested graph conditions have been shown to be equivalent to first-order logic over graphs [HP09, Ren04]. The satisfiability problem for first-order logic being undecidable thus makes a direct computation of matches for which NGC satisfaction somehow depends on the relevant subgraph impossible in the general case.
In order to still enable a localized tracking of changes to the results of an extended graph query, we therefore employ the following overapproximation for the set of matches where NGC satisfaction may change as a result of local modifications:
@@
  • $\psi=\exists(a:Q\rightarrow Q^{\prime},\psi^{\prime})$ and there exists a graph morphism $m^{\prime}:Q^{\prime}\rightarrow H$ such that $m=m^{\prime}\circ a$ and $m^{\prime}(Q^{\prime})\cap H_{p}\neq\emptyset\text{ or }m^{\prime}\text{ is subgraph satisfaction dependent}$
@@
We furthermore define a subgraph-restricted, that is, local graph modification based on the basic notion of graph modifications introduced in Section 2.1. In a subgraph-restricted graph modification, changes are limited to relevant subgraphs of source and target graph and relevant subgraphs are mapped onto each other:
@@
The relationship between the graphs in the above definition is visualized in the diagram in Figure 10. $H$ is the source graph of the modification, which is modified into $H^{\prime}$ via the intermediate graph $K$. $H_{p}$ and $H^{\prime}_{p}$ are the relevant subgraphs of $H$ respectively $H^{\prime}$, which have to be mapped onto each other via $f\circ g^{-1}$ and $g\circ f^{-1}$. $H_{s}$ and $H^{\prime}_{s}$ constitute the remainder of $H$ and $H^{\prime}$ and are required to be isomorphic, which means that the subgraph-restricted graph modification cannot make changes outside the relevant subgraph. Note that $H_{s}$ and $H^{\prime}_{s}$ may share nodes with $H_{p}$ respectively $H^{\prime}_{p}$ if these nodes are adjacent to edges in $H\setminus H_{p}$ respectively $H^{\prime}\setminus H^{\prime}_{p}$ to avoid dangling edges.
[…]
Figure 10. Relationship between graphs in Definition 4.2
@@
  • For a nested graph condition of the form $\psi=\exists(a:Q\rightarrow Q^{\prime},\psi^{\prime})$, $(N^{sat},p^{sat})=localize^{sat}(Q,\psi)$ consists of the localized RETE net for the plain pattern $Q$, $(N^{\Phi}_{Q},p^{\Phi}_{Q})=localize(Q)$, the localized RETE net for the plain pattern $Q^{\prime}$, $(N^{\Phi}_{Q^{\prime}},p^{\Phi}_{Q^{\prime}})=localize(Q^{\prime})$, the RETE net $(N^{sat}_{(Q^{\prime},\psi^{\prime})},p^{sat}_{(Q^{\prime},\psi^{\prime})})=localize^{sat}(Q^{\prime},\psi^{\prime})$, a marking-sensitive union node $[\cup]^{\Phi}$ with dependencies $p^{\Phi}_{Q^{\prime}}$ and $p^{sat}_{(Q^{\prime},\psi^{\prime})}$, a request projection structure given by $RPS^{\infty}_{r}=RPS^{\infty}([\cup]^{\Phi},N^{\Phi}_{Q})$, and a marking-sensitive semi-join node $p^{sat}=[\ltimes]^{\Phi}$ along $a$ with left dependency $p^{\Phi}_{Q}$ and right dependency $[\cup]^{\Phi}$.
@@
A localized RETE net for computing subgraph-satisfaction dependent matches for the example query in Figure 3 is displayed in Figure 12 alongside a consistent configuration for the host graph in Figure 13. It consists of the localized RETE net in Figure 5 and the structure created for handling the nested graph condition according to Figure 11.
Notably, there exist no matches for the base pattern $Q$ that touch the relevant subgraph, which only consists of the interface node $i_{1}$. Instead, the computation of matches for $Q$ is controlled via the request projection structure that feeds matches for the right dependency of the semi-join node into the localized RETE net for $Q$.
Compared to Figure 8, propagation of matches is essentially reversed: First, the local navigation structure $LNS^{c\rightarrow i}$ extracts all matches for the pattern $Q^{\prime}$ that touch the relevant subgraph. The request projection structure connected to the union node then causes the local navigation structure $LNS^{c\rightarrow f}$ to extract the match $m_{1.2}$. The leftmost request projection structure finally ensures that the required complementary match $m_{1.1}$ is extracted by the local navigation structure $LNS^{p\rightarrow c}$.
[…]
Figure 12. Localized RETE net for computing subgraph-satisfaction dependent matches for the example query in Figure 3 with a consistent configuration for the host graph in Figure 13
r1:Projp1:Pkgc1:Classf1:Fieldi1:Intfpececefeiep2:Pkgc2:Classf2:Fieldpecefe m1 m1.1 m1.1.1 m1.2 m1.2.1 m2 m1.3
Figure 13. Example host graph with relevant subgraph marked in bold
A consistent configuration for the resulting RETE net then indeed guarantees a complete query result with respect to Definition 4.2:
@@
(Consistent configurations for RETE nets localized via $localize^{sat}$ yield all subgraph satisfaction dependent matches.) Let $H$ be a graph, $H_{p}\subseteq H$, $(Q,\psi)$ an extended graph query, and $\mathcal{C}^{\Phi}$ a consistent configuration for the RETE net $(N^{sat},p^{sat})=localize^{sat}(Q,\psi)$. It then holds that $\forall m\in\mathcal{M}^{Q}_{H}:m\text{ is subgraph satisfaction dependent }\Rightarrow m\in\mathcal{R}^{\Phi}_{X}(p^{sat},\mathcal{C}^{\Phi})$.
@@
  • For a nested graph condition of the form $\psi=\exists(a:Q\rightarrow Q^{\prime},\psi^{\prime})$, the result of $order^{sat}$ is given by $order^{sat}(N^{sat})=order(N^{\Phi}_{Q^{\prime}})\circ order^{sat}(N^{sat}_{(Q^{\prime},\psi^{\prime})})\circ[\cup]^{\Phi}\circ toposort(RPS^{\infty}_{r})^{-1}\circ order(N^{\Phi}_{Q})\circ[\ltimes]^{\Phi}$.
@@
(Execution of localized RETE nets via $order^{sat}$ yields consistent configurations.) Let $H$ be a graph, $H_{p}\subseteq H$, $(Q,\psi)$ a graph query, and $\mathcal{C}^{\Phi}_{0}$ an arbitrary starting configuration for the marking-sensitive RETE net $(N^{sat},p^{sat})=localize^{sat}(Q,\psi)$. Executing $(N^{sat},p^{sat})$ via $O=order^{sat}(N^{sat})$ then yields a consistent configuration $\mathcal{C}^{\Phi}=execute(O,N^{sat},H,H_{p},\mathcal{C}^{\Phi}_{0})$.
@@
(Execution of RETE nets localized via $localize^{sat}$ yields all subgraph satisfaction dependent matches.) Let $H$ be a graph, $H_{p}\subseteq H$, and $(Q,\psi)$ a graph query. Furthermore, let $\mathcal{C}^{\Phi}_{0}$ be an arbitrary starting configuration for the marking-sensitive RETE net $(N^{sat},p^{sat})=localize^{sat}(Q,\psi)$ and $\mathcal{C}^{\Phi}=execute(order^{sat}(N^{sat}),N^{sat},H,H_{p},\mathcal{C}^{\Phi}_{0})$. It then holds that $\forall m\in\mathcal{M}^{Q}_{H}:m\text{ is subgraph satisfaction dependent }\Rightarrow m\in\mathcal{R}^{\Phi}_{X}(p^{sat},\mathcal{C}^{\Phi})$.
@@
The result $(N^{\Delta},p^{\Delta})=localize^{\Delta}(Q,\psi)$ of this construction via the procedure $localize^{\Delta}$ is visualized in Figure 14. The resulting RETE net consists of the localized version $(N^{\Phi}_{Q},p^{\Phi}_{Q})=localize(N^{Q},p^{Q})$ of a regular RETE net $(N^{Q},p^{Q})$ with height $h$ for the plain graph query $Q$, a marking filter node $[\phi>h]^{\Phi}$ with dependency $p^{\Phi}_{Q}$, the RETE net $(N^{sat}_{(Q,\psi)},p^{sat}_{(Q,\psi)})=localize^{sat}(Q,\psi)$, and a marking-sensitive union node $[\cup]^{\Phi}$ with dependencies $[\phi>h]^{\Phi}$ and $p^{sat}_{(Q,\psi)}$. In addition, $N^{\Delta}$ comprises the RETE net $(N^{\Phi}_{(Q,\psi)},p^{\Phi}_{(Q,\psi)})=localize^{\Psi}(Q,\psi)$, a request projection structure $RPS^{\infty}_{l}=RPS^{\infty}([\cup]^{\Phi},N^{\Phi}_{(Q,\psi)})$, and a marking-sensitive semi-join $[\ltimes]^{\Phi}$ with left dependency $[\cup]^{\Phi}$ and right dependency $p^{\Phi}_{(Q,\psi)}$.
Additionally, another copy of the RETE net for computing subgraph satisfaction dependent matches, $(N^{sat^{\prime}}_{(Q,\psi)},p^{sat^{\prime}}_{(Q,\psi)})=localize^{sat}(Q,\psi)$, is created for computing subgraph satisfaction dependent matches in $H^{\prime}$. These matches then have to be translated into matches into $H$ along the graph morphisms $f$ and $g$ via a dedicated marking-sensitive translation node $[\xrightarrow{f\circ g^{-1}}]^{\Phi}$ with dependency $p^{sat^{\prime}}_{(Q,\psi)}$, which is also added to $N^{\Delta}$. The target result set of this marking-sensitive translation node is given by $\mathcal{R}^{\Phi}([\xrightarrow{f\circ g^{-1}}]^{\Phi},N^{\Delta},H,H_{p},\mathcal{C}^{\Phi})=\{(m,\phi)|m\in\mathcal{M}^{Q}_{H}\wedge\exists(m^{\prime},\phi)\in\mathcal{C}^{\Phi}(p^{sat^{\prime}}):f\circ g^{-1}\circ m^{\prime}=m\}$. The node $[\xrightarrow{f\circ g^{-1}}]^{\Phi}$ can then be added as an additional dependency to $[\cup]^{\Phi}$.
@@
An execution order for a RETE net $(N^{\Delta},p^{\Delta})=localize^{\Delta}(Q,\psi)$ can simply be constructed as $order^{\Delta}(N^{\Delta})=order(N^{\Phi}_{Q})\circ[\phi>h]^{\Phi}\circ order^{sat}(N^{sat}_{(Q,\psi)})\circ[\xrightarrow{f\circ g^{-1}}]^{\Phi}\circ[\cup]^{\Phi}\circ toposort(RPS^{\infty}_{l})^{-1}\circ order^{\Psi}(N^{\Phi}_{(Q,\psi)})\circ[\ltimes]^{\Phi}$:
@@
(Execution of localized RETE nets via $order^{\Delta}$ yields consistent configurations.) Let $(Q,\psi)$ be an extended graph query and $H$ a host graph with relevant subgraph $H_{p}$ and $\mathcal{C}^{\Phi}_{0}$ an arbitrary starting configuration for the marking-sensitive RETE net $(N^{\Delta},p^{\Delta})=localize^{\Delta}(Q,\psi)$. Executing $(N^{\Delta},p^{\Delta})$ via $O=order^{\Delta}(N^{\Delta})$ then yields a consistent configuration $\mathcal{C}^{\Phi}=execute(O,N^{\Delta},H,H_{p},\mathcal{C}^{\Phi}_{0})$.
@@ 5. Evaluation
We aim to investigate whether RETE net localization can improve performance of query execution in scenarios where the relevant subgraph constitutes only a fraction of the full model, considering initial query execution time, execution time for incrementally processing model updates, and memory consumption as performance indicators. We experiment [^7] Experiments were executed on a Linux SMP Debian 4.19.67-2 computer with Intel Xeon E5-2630 CPU (2.3 GHz clock rate) and 386 GB system memory running OpenJDK 11.0.6. Reported measurements correspond to the arithmetic mean of measurements for 10 runs. Memory measurements were obtained using the Java Runtime class. To represent graph data, all experiments use EMF [emf] object graphs that enable reverse navigation of edges. Our implementation is available under [BG24c]. More details and query visualizations can be found in [BG24b] and [BG24c] or in Appendix B. with the following querying techniques:
@@ Appendix A Technical Details
Theorem 3.6.
(Matches are marked $\infty$ in query results of RETE nets localized via $localize$ iff they touch the relevant subgraph.) Let $H$ be a graph, $H_{p}\subseteq H$, $(N,p)$ a well-formed RETE net, and $Q=query(p)$. Furthermore, let $\mathcal{C}^{\Phi}$ be a consistent configuration for the localized RETE net $(N^{\Phi},p^{\Phi})=localize(N,p)$. It then holds that $\forall(m,\phi)\in\mathcal{C}^{\Phi}(p^{\Phi}):m(Q)\cap H_{p}\neq\emptyset\Leftrightarrow\phi=\infty$.
@@
Theorem 3.10.
(RETE net localization via $localize$ introduces only a constant factor overhead on effective configuration size.) Let $H$ be an edge-dominated graph, $H_{p}\subseteq H$, $(N,p)$ a well-formed RETE net with $Q=query(p)$, $\mathcal{C}$ a consistent configuration for $(N,p)$ for host graph $H$, and $\mathcal{C}^{\Phi}$ a consistent configuration for the marking-sensitive RETE net $(N^{\Phi},p^{\Phi})=localize(N,p)$ for host graph $H$ and relevant subgraph $H_{p}$. It then holds that $\sum_{n^{\Phi}\in V^{N^{\Phi}}}\sum_{(m,\phi)\in\mathcal{C}(n^{\Phi})}|m|\leq 7\cdot|\mathcal{C}|_{e}$.
@@
Theorem 4.3.
(Consistent configurations for RETE nets localized via $localize^{\Psi}$ yield correct query results under the relevant subgraph.) Let $H$ be a graph, $H_{p}\subseteq H$, and $(Q,\psi)$ an extended graph query. Furthermore, let $\mathcal{C}^{\Phi}$ be a consistent configuration for the localized RETE net $(N^{\Phi},p^{\Phi})=localize^{\Psi}(Q,\psi)$. The set of matches given by the stripped result set $\mathcal{R}^{\Phi}_{X}(p^{\Phi},\mathcal{C}^{\Phi})$ is then correct under $H_{p}$.
@@
Theorem 4.4.
(Execution of localized RETE nets via $order^{\Psi}$ yields consistent configurations.) Let $H$ be a graph, $H_{p}\subseteq H$, $(Q,\psi)$ a graph query, and $\mathcal{C}^{\Phi}_{0}$ an arbitrary starting configuration for the marking-sensitive RETE net $(N^{\Phi},p^{\Phi})=localize^{\Psi}(Q,\psi)$. Executing $(N^{\Phi},p^{\Phi})$ via $O=order^{\Psi}(N^{\Phi})$ then yields a consistent configuration $\mathcal{C}^{\Phi}=execute(O,N^{\Phi},H,H_{p},\mathcal{C}^{\Phi}_{0})$.
@@
Theorem 4.6.
(RETE net localization via $localize^{\Psi}$ introduces only a constant factor overhead on effective configuration size.) Let $H$ be an edge-dominated graph, $H_{p}\subseteq H$, $(N,p)$ a RETE net created via the procedure described in [BG23] for the extended graph query $(Q,\psi)$, $\mathcal{C}$ a consistent configuration for $(N,p)$ for host graph $H$, and $\mathcal{C}^{\Phi}$ a consistent configuration for the marking-sensitive RETE net $(N^{\Phi},p^{\Phi})=localize^{\Psi}(Q,\psi)$ corresponding to $(N,p)$ for host graph $H$ and relevant subgraph $H_{p}$. It then holds that $\sum_{n^{\Phi}\in V^{N^{\Phi}}}\sum_{(m,\phi)\in\mathcal{C}^{\Phi}(n^{\Phi})}|m|\leq 7\cdot|\mathcal{C}|_{e}$.
@@
For a nested condition of the form $\psi=\exists(a:Q\rightarrow Q^{\prime},\psi^{\prime})$, $N^{\Phi}$ consists of the localized RETE net for $Q$, $(N^{\Phi}_{Q},p^{\Phi}_{Q})=localize(Q)$, the RETE net $(N^{\Phi}_{(Q^{\prime},\psi^{\prime})},p^{\Phi}_{(Q^{\prime},\psi^{\prime})})=localize^{\Psi}(Q^{\prime},\psi^{\prime})$, a request projection structure $RPS^{\infty}_{l}=RPS^{\infty}(p^{\Phi}_{Q},N^{\Phi}_{(Q^{\prime},\psi^{\prime})})$, and a marking-sensitive semi-join node $p^{\Phi}=[\ltimes]^{\Phi}$.